HYPERFINE STRUCTURE OF THE B ^{2}Σ^{+} − X ^{2}Σ^{+} TRANSITION OF LaS

\underline{A. S-C. CHEUNG}, W. S. TAM, J. W-H. LEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; SHENGGUI HE, Department of Chemical Physics, University of Science and Technology of China, Hefei, Anhui, China.